You discover a novel unicellular, sexually reproducing organism. The ‘A’ (or ‘a’) locus is 38 map units away from the ‘B’ (or ‘b’) locus. Suppose that ‘A’ encodes cilia that help the cell to move (and is dominant to ‘a’ which is incapable of moving) and that ‘B’ encodes ability to perform photosynthesis (and is dominant to ‘b’ which is incapable of photosynthesis).

  1. If you take the offspring of this cross and mate one that can move but can’t perform photosynthesis with one that can move and photosynthesize, what fraction of offspring of each phenotype do you expect? Assume that the loci are still 38 map units away from each other.
Answer #1

Recombination frequency is directly proportional to map distance

Recombinant will be 38%

Half of the parents = 100- 38= 62, ( 31 each parental type,)

Recombinant = 38 ( 17 of each recombinant)

Explanation has been attached below